I am trying to help my duahgter with her pre-algebra. Her paper only states the following: y = 2x-5 x-y = 2
She does not know the formula to figure this out. I have forgotten...can you please help us?

Answer by nerdybill(7384)

You can apply the "substitution method".
.
y = 2x-5 (equation 1)
x-y = 2 (equation 2)
.
Use the definition of 'y' from equation 1 and substitute it into equation 2 -- then solve for x:
x-y = 2
Substitute our definition of y:
x-(2x-5) = 2
x-2x+5 = 2
-x + 5 = 2
5 = x + 2
3 = x (solution for x)
.
Substitute the above into equation 2 and solve for y:
x-y = 2
3-y = 2
3 = y+2
1 = y (solution of y)
.
Final answer:
(x,y) = (3,1)